OpenStudy (anonymous):

whats the midpoint of a line segment whose endpoints are (0,1/2) and (0,3/4)

OpenStudy (alekos):

midpoint = [(x1+x2)/2,(y1+y2)/2]

OpenStudy (anonymous):

Thats the formula right??

OpenStudy (alekos):

yes it is

OpenStudy (alekos):

x1=x2=0 and y1=1/2, y2=3/4

OpenStudy (anonymous):

so do i just put the numbers in the equation then solve it ?

OpenStudy (alekos):

yes you do

OpenStudy (anonymous):

ok thanks

OpenStudy (alekos):

yw

OpenStudy (anonymous):

ok so this is what i got now! [(0)/2, (5/4)/2? Is that right

OpenStudy (anonymous):

is anybody there

OpenStudy (phi):

yes that is correct, but can be simplified. 0/2 is 0 \[ \frac{\frac{5}{4}}{2} = \frac{5}{4} \cdot \frac{1}{2} = \frac{5}{8} \] so the final answer is (0, 5/8)